# Assimilative Causal Inference

### v1

@@ 5.2 A Stochastic Model Capturing the El Niño-Southern Oscillation (ENSO) Diversity
El Niño-Southern Oscillation (ENSO) is a dominant climate phenomenon characterized by quasi-regular periodic behaviors in sea surface temperatures (SSTs) and in the atmospheric circulation across the equatorial Pacific Ocean [10, 39]. ENSO exhibits remarkable diversity in its spatial patterns, temporal evolution, and impacts, which can be broadly categorized into two main types: Eastern Pacific (EP) and Central Pacific (CP) El Niños, where an anomalous warming center occurs in the eastern and central Pacific, respectively [6]. The opposite phases with anomalous cooling SSTs are called La Niña. Understanding the causes of different ENSO events is crucial for improving climate predictions and mitigating socio-economic consequences.
Although few models can accurately capture ENSO diversity, a recently developed stochastic conceptual model successfully reproduces its diverse behaviors and non-Gaussian statistics [8]. This model has been highlighted in a recent review [47], making it a suitable testbed for studying El Niño diversity. Mathematical details are provided in the Supporting Information. The model consists of six state variables: ocean zonal current in the CP ($u$), western Pacific (WP) thermocline depth ($h_{W}$), CP SST ($T_{C}$), EP SST ($T_{E}$), atmospheric winds ($\tau$, intraseasonal), and decadal variation ($I$, decadal). The first four variables ($u$, $h_{W}$, $T_{C}$, and $T_{E}$) operate on interannual timescales. As a nonlinear system with state-dependent noise, the model generates extreme events and intermittency. The two SST variables ($T_{C}$ and $T_{E}$) allow reconstruction of spatiotemporal SST patterns across the equatorial Pacific, providing an intuitive way to identify different ENSO event types.
@@
Figure 5.2: ACI values and CIRs of the stochastic conceptual model for the ENSO diversity with $T_{E}$ as the effect variable over a six-year model-simulated period (where EEP EN and MEP EN stand for extreme and moderate EP El Niños, respectively). Panel (a): Hovmöller diagram of the SST anomalies in the equatorial Pacific reconstructed from $T_{C}$ and $T_{E}$ series via a spatiotemporal multivariate regression technique [8]. Panel (b): ACI values from $T_{C}$ to $T_{E}$ as a function of time. Panel (c): Time series of the observable $T_{C}$ and the objective CIR represented by a whisker plot emanating forward in time from the value of $T_{C}$ at each month over this six-year period. Panels (d) and (e): Same as Panels (b) and (c) but for the causal link from $h_{W}$ to $T_{E}$. Panels (f) and (g): Same as Panels (b) and (c) but for the causal link from $\tau$ to $T_{E}$.
@@ 6 Conclusion and Discussions
In this paper, the assimilative causal inference (ACI) framework is developed for the detection of instantaneous causal relationships and of their associated causal influence range (CIR) in complex dynamical systems. ACI leverages Bayesian data assimilation to trace causes backward from observed effects. It uniquely identifies dynamic causal interactions without requiring observations from candidate causes, accommodates short datasets, and scales efficiently to high dimensions. Crucially, it provides online tracking of causal roles, which may reverse intermittently, and facilitates a mathematically rigorous criterion for the causal influence range, revealing how far effects propagate temporally. Its ability to detect causal influence ranges and handle high-dimensional systems highlights its potential for applications in climate science, neuroscience, and other fields. Numerical tests on nonlinear systems with extreme events and regime transitions demonstrate the effectiveness and robustness of the framework.
Future work will focus on several important directions. First, we will study the impact of model error on causal inference, as model error is common in practice. Understanding how model error affects causal relationships may also provide opportunities to detect and correct model errors using observational data. Second, we will extend the framework to attribute observed regime transitions and other dynamic phenomena, complementing the current forward-in-time CIR analysis with backward-in-time event attribution. Finally, we will develop efficient computational algorithms based on ensemble data assimilation, thus advancing the application of the ACI to highly complicated systems.
@@
No external datasets were used in this research; all data in the simulations are model-generated.
@@
  • [1] M. Andreou, N. Chen, and Y. Li (2024) An adaptive online smoother with closed-form solutions and information-theoretic lag selection for conditional gaussian nonlinear systems. arXiv preprint arXiv:2411.05870. Cited by: §3.1.
  • [2] J. D. Angrist and J. Pischke (2009) Mostly harmless econometrics: an empiricist’s companion. Princeton university press. Cited by: §1.
  • [3] L. Barnett, A. B. Barrett, and A. K. Seth (2009) Granger causality and transfer entropy are equivalent for Gaussian variables. Physical review letters 103 (23), pp. 238701. Cited by: §1, §2.1.
  • [4] J. E. Brand, X. Zhou, and Y. Xie (2023) Recent developments in causal inference and machine learning. Annual Review of Sociology 49 (1), pp. 81–110. Cited by: §1.
  • [5] C. Cafaro, W. M. Lord, J. Sun, and E. M. Bollt (2015) Causation entropy from symbolic representations of dynamical systems. Chaos: An interdisciplinary journal of nonlinear science 25 (4). Cited by: §1.
  • [6] A. Capotondi, A. T. Wittenberg, M. Newman, E. Di Lorenzo, J. Yu, P. Braconnot, J. Cole, B. Dewitte, B. Giese, E. Guilyardi, et al. (2015) Understanding ENSO diversity. Bulletin of the American Meteorological Society 96 (6), pp. 921–938. Cited by: §5.2.
  • [7] A. Carrassi, M. Bocquet, A. Hannart, and M. Ghil (2017) Estimating model evidence using data assimilation. Quarterly Journal of the Royal Meteorological Society 143 (703), pp. 866–880. Cited by: §1.
  • [8] N. Chen, X. Fang, and J. Yu (2022) A multiscale model for El Niño complexity. npj Climate and Atmospheric Science 5 (1), pp. 16. Cited by: Figure 5.2, §5.2.
  • [9] N. Chen (2023) Stochastic methods for modeling and predicting complex dynamical systems. Springer. Cited by: §2.3.
  • [10] A. J. Clarke (2008) An introduction to the dynamics of El Niño and the southern oscillation. Elsevier. Cited by: §5.2.
  • [11] T. M. Cover (1999) Elements of information theory. John Wiley & Sons. Cited by: §1.
  • [12] G. Evensen, F. C. Vossepoel, and P. J. Van Leeuwen (2022) Data assimilation fundamentals: a unified formulation of the state and parameter estimation problem. Springer Nature. Cited by: §2.2.
  • [13] F. Falasca, P. Perezhogin, and L. Zanna (2023) Causal inference in spatiotemporal climate fields through linear response theory. arXiv preprint arXiv:2306.14433. Cited by: §1.
  • [14] A. Fowler and P. Jan Van Leeuwen (2013) Observation impact in data assimilation: the effect of non-Gaussian observation error. Tellus A: Dynamic Meteorology and Oceanography 65 (1), pp. 20035. Cited by: §2.3.
  • [15] S. Frenzel and B. Pompe (2007) Partial mutual information for coupling analysis of multivariate time series. Physical review letters 99 (20), pp. 204101. Cited by: §1.
  • [16] C. W. Granger (1969) Investigating causal relations by econometric models and cross-spectral methods. Econometrica: journal of the Econometric Society, pp. 424–438. Cited by: §1, §2.1.
  • [17] A. Hannart, A. Carrassi, M. Bocquet, M. Ghil, P. Naveau, M. Pulido, J. Ruiz, and P. Tandeo (2016) DADA: data assimilation for the detection and attribution of weather and climate-related events. Climatic Change 136, pp. 155–174. Cited by: §1.
  • [18] A. Hannart, J. Pearl, F. Otto, P. Naveau, and M. Ghil (2016) Causal counterfactual theory for the attribution of weather and climate-related events. Bulletin of the American Meteorological Society 97 (1), pp. 99–110. Cited by: §1.
  • [19] F. Jin (1997) An equatorial ocean recharge paradigm for ENSO. Part I: Conceptual model. Journal of the atmospheric sciences 54 (7), pp. 811–829. Cited by: §5.2.
  • [20] M. Kocaoglu, K. Shanmugam, and E. Bareinboim (2017) Experimental design for learning causal graphs with latent variables. Advances in Neural Information Processing Systems 30. Cited by: §1.
  • [21] K. Law, A. Stuart, and K. Zygalakis (2015) Data assimilation. Cham, Switzerland: Springer 214, pp. 52. Cited by: §2.2.
  • [22] X. S. Liang and R. Kleeman (2005) Information transfer between dynamical system components. Physical review letters 95 (24), pp. 244101. Cited by: §1, §1, §1.
  • [23] V. Lucarini and M. D. Chekroun (2024) Detecting and attributing change in climate and complex systems: Foundations, Green’s functions, and nonlinear fingerprints. Physical Review Letters 133 (24), pp. 244201. Cited by: §1.
  • [24] A. J. Majda and N. Chen (2018) Model error, information barriers, state estimation and prediction in complex multiscale systems. Entropy 20 (9), pp. 644. Cited by: §2.3.
  • [25] A. J. Majda and J. Harlim (2012) Physics constrained nonlinear regression models for time series. Nonlinearity 26 (1), pp. 201. Cited by: §5.1.
  • [26] D. Mønster, R. Fusaroli, K. Tylén, A. Roepstorff, and J. F. Sherson (2017) Causal inference from noisy time-series data—testing the convergent cross-mapping algorithm in the presence of noise and external influence. Future Generation Computer Systems 73, pp. 52–62. Cited by: §1.
  • [27] S. Morgan (2015) Counterfactuals and causal inference. Cambridge University Press. Cited by: §1.
  • [28] J. Pearl et al. (2000) Models, reasoning and inference. Cambridge, UK: CambridgeUniversityPress 19 (2), pp. 3. Cited by: §1.
  • [29] J. Peters, D. Janzing, and B. Schölkopf (2017) Elements of causal inference: foundations and learning algorithms. The MIT Press. Cited by: §1.
  • [30] M. Prosperi, Y. Guo, M. Sperrin, J. S. Koopman, J. S. Min, X. He, S. Rich, M. Wang, I. E. Buchan, and J. Bian (2020) Causal inference and counterfactual prediction in machine learning for actionable healthcare. Nature Machine Intelligence 2 (7), pp. 369–375. Cited by: §1.
  • [31] A. Radebach, R. V. Donner, J. Runge, J. F. Donges, and J. Kurths (2013) Disentangling different types of El Niño episodes by evolving climate network analysis. Physical Review E—Statistical, Nonlinear, and Soft Matter Physics 88 (5), pp. 052807. Cited by: §1.
  • [32] S. Reich and C. Cotter (2015) Probabilistic forecasting and bayesian data assimilation. Cambridge University Press. Cited by: §2.2.
  • [33] D. B. Rubin (1974) Estimating causal effects of treatments in randomized and nonrandomized studies.. Journal of educational Psychology 66 (5), pp. 688. Cited by: §1.
  • [34] J. Runge, S. Bathiany, E. Bollt, G. Camps-Valls, D. Coumou, E. Deyle, C. Glymour, M. Kretschmer, M. D. Mahecha, J. Muñoz-Marí, et al. (2019) Inferring causation from time series in earth system sciences. Nature communications 10 (1), pp. 2553. Cited by: §1.
  • [35] J. Runge, J. Heitzig, N. Marwan, and J. Kurths (2012) Quantifying causal coupling strength: A lag-specific measure for multivariate time series related to transfer entropy. Physical Review E—Statistical, Nonlinear, and Soft Matter Physics 86 (6), pp. 061121. Cited by: §1.
  • [36] J. Runge, P. Nowack, M. Kretschmer, S. Flaxman, and D. Sejdinovic (2019) Detecting and quantifying causal associations in large nonlinear time series datasets. Science advances 5 (11), pp. eaau4996. Cited by: §1.
  • [37] J. Runge, V. Petoukhov, J. F. Donges, J. Hlinka, N. Jajcay, M. Vejmelka, D. Hartman, N. Marwan, M. Paluš, and J. Kurths (2015) Identifying causal gateways and mediators in complex spatio-temporal systems. Nature communications 6 (1), pp. 8502. Cited by: §1.
  • [38] A. Rupe, D. DeSantis, C. Bakker, P. Kooloth, and J. Lu (2024) Causal discovery in nonlinear dynamical systems using Koopman operators. arXiv preprint arXiv:2410.10103. Cited by: §1.
  • [39] E. S. Sarachik and M. A. Cane (2010) The El Nino-southern oscillation phenomenon. Cambridge University Press. Cited by: §5.2.
  • [40] S. Särkkä and L. Svensson (2023) Bayesian filtering and smoothing. Vol. 17, Cambridge university press. Cited by: §2.2.
  • [41] T. Schreiber (2000) Measuring information transfer. Physical review letters 85 (2), pp. 461. Cited by: §1, §2.1.
  • [42] K. Shanmugam, M. Kocaoglu, A. G. Dimakis, and S. Vishwanath (2015) Learning causal graphs with small interventions. Advances in Neural Information Processing Systems 28. Cited by: §1.
  • [43] G. Sugihara, R. May, H. Ye, C. Hsieh, E. Deyle, M. Fogarty, and S. Munch (2012) Detecting causality in complex ecosystems. science 338 (6106), pp. 496–500. Cited by: §1.
  • [44] J. Sun and E. M. Bollt (2014) Causation entropy identifies indirect influences, dominance of neighbors and anticipatory couplings. Physica D: Nonlinear Phenomena 267, pp. 49–57. Cited by: §1.
  • [45] J. Sun, D. Taylor, and E. M. Bollt (2015) Causal network inference by optimal causation entropy. SIAM Journal on Applied Dynamical Systems 14 (1), pp. 73–106. Cited by: §1.
  • [46] S. Varambally, Y. Ma, and R. Yu (2023) Discovering mixtures of structural causal models from time series data. arXiv preprint arXiv:2310.06312. Cited by: §1.
  • [47] J. Vialard, F. Jin, M. McPhaden, A. Fedorov, W. Cai, S. An, D. Dommenget, X. Fang, M. Stuecker, C. Wang, et al. (2025) The El Niño Southern Oscillation (ENSO) recharge oscillator conceptual model: Achievements and future prospects. Reviews of Geophysics 63 (1), pp. e2024RG000843. Cited by: §5.2.
  • [48] Q. Xu (2007) Measuring information content from observations for data assimilation: Relative entropy versus Shannon entropy difference. Tellus A: Dynamic Meteorology and Oceanography 59 (2), pp. 198–209. Cited by: §2.3.
  • [49] X. Yang, Z. Wang, C. Wang, and Y. Lai (2023) Finding causal gateways of precipitation over the contiguous United States. Geophysical Research Letters 50 (4), pp. e2022GL101942. Cited by: §1.
  • [50] H. Ye, E. R. Deyle, L. J. Gilarranz, and G. Sugihara (2015) Distinguishing time-delayed causal interactions using convergent cross mapping. Scientific reports 5 (1), pp. 14750. Cited by: §1.
Supporting Information
@@ A.2.1 Underlying Dynamics
We assume that the evolution of the partially observed process $(\mathbf{x},\mathbf{y})$ is governed by the following stochastic system of coupled Itô diffusions over $t\in[0,T]$ [21]:
@@
where $\mathbf{W}_{1}\in\mathbb{R}^{d_{1}}$ and $\mathbf{W}_{2}\in\mathbb{R}^{d_{2}}$ are two real-valued independent Wiener processes that are also independent from the distribution of $(\mathbf{x}(0),\mathbf{y}(0))$. In (1.1), we assume the following [21, 9, 19]. (I) Almost every sample path of $\mathbf{x}$ and $\mathbf{y}$ is continuous in $[0,T]$. (II) The functions appearing in (1.1) are predictable for each $\mathbf{x}\in\mathbb{R}^{k}$ and $\mathbf{y}\in\mathbb{R}^{l}$. (III) Sufficient regularity conditions are enforced such that (1.1) has a unique solution. (IV) The diffusion coefficients of $\mathbf{x}$ do not depend on $\mathbf{y}$: $\boldsymbol{\Sigma}^{\mathbf{x}}_{m}(t,\mathbf{x},\mathbf{y})=\boldsymbol{\Sigma}^{\mathbf{x}}_{m}(t,\mathbf{x})$, for $m=1,2$. (V) The sum of the row-based Gramians of the observational noise feedback matrices, $(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})(t,\mathbf{x}):=\boldsymbol{\Sigma}^{\mathbf{x}}_{1}(t,\mathbf{x})\boldsymbol{\Sigma}^{\mathbf{x}}_{1}(t,\mathbf{x})^{\mathtt{T}}+\boldsymbol{\Sigma}^{\mathbf{x}}_{2}(t,\mathbf{x})\boldsymbol{\Sigma}^{\mathbf{x}}_{2}(t,\mathbf{x})^{\mathtt{T}}$, is invertible for each $t\in[0,T]$ and $\mathbf{x}\in\mathbb{R}^{k}$.
(IV)–(V) function as identifiability conditions, since they ensure that the conditional distribution of $\mathbf{y}$ given the data of $\mathbf{x}$ contains all available information about $\mathbf{y}$ in (1.1) [21].

### v2

@@ 5.2 A Stochastic Model Capturing the El Niño-Southern Oscillation (ENSO) Diversity
El Niño–Southern Oscillation (ENSO) is a dominant climate phenomenon characterized by quasi-regular periodic behaviors in sea surface temperatures (SSTs) and in the atmospheric circulation across the equatorial Pacific Ocean [clarke2008introduction, sarachik2010nino]. ENSO exhibits remarkable diversity in its spatial patterns, temporal evolution, and impacts, which can be broadly categorized into two main types: Eastern Pacific (EP) and Central Pacific (CP) El Niños, where an anomalous warming center occurs in the eastern and central Pacific, respectively [capotondi2015understanding]. The opposite phases with anomalous cooling SSTs are called La Niña. Understanding the causes of different ENSO events is crucial for improving climate predictions and mitigating socio-economic consequences.
Although few models can accurately capture ENSO diversity, a recently developed stochastic conceptual model successfully reproduces its diverse behaviors and non-Gaussian statistics [chen2022multiscale]. This model has been highlighted in a recent review [vialard2025nino], making it a suitable testbed for studying the ENSO diversity. Mathematical details are provided in the Supplementary Information. The model consists of six state variables: ocean zonal current in the CP ($u$), western Pacific (WP) thermocline depth ($h_{W}$), CP SST ($T_{C}$), EP SST ($T_{E}$), atmospheric winds ($\tau$, intraseasonal), and background Walker circulation ($I$, decadal). The first four variables ($u$, $h_{W}$, $T_{C}$, and $T_{E}$) operate on interannual timescales. As a nonlinear system with state-dependent noise, the model generates extreme events and intermittency. The two SST variables allow reconstruction of spatiotemporal SST patterns across the equatorial Pacific, providing an intuitive way to identify different ENSO event types.
@@
Figure 3: Conditional ACI values and CIRs of the stochastic conceptual model for the ENSO diversity with $T_{E}$ as the effect variable over a six-year model-simulated period (where EEP EN and MEP EN stand for extreme and moderate EP El Niños, respectively, and LN for La Niña). Panel (a): Hovmöller diagram of the SST anomalies in the equatorial Pacific reconstructed from $T_{C}$ and $T_{E}$ via a spatiotemporal multivariate regression technique [chen2022multiscale]. Panel (b): ACI values from $T_{C}$ to $T_{E}$ as a function of time. Panel (c): Time series of the observable $T_{C}$ and the objective CIR represented by a whisker plot emanating forward in time from the value of $T_{C}$ at each month over this six-year period. Panels (d) and (e): Same as Panels (b) and (c) but for the causal link from $h_{W}$ to $T_{E}$. Panels (f) and (g): Same as Panels (b) and (c) but for the causal link from $\tau$ to $T_{E}$.
@@ 6 Conclusion and Discussions
In this paper, the assimilative causal inference (ACI) framework is developed for the detection of instantaneous causal relationships and of their associated causal influence range (CIR) in complex dynamical systems. ACI leverages Bayesian data assimilation to trace causes backward from observed effects. It uniquely identifies dynamic causal interactions without requiring observations from candidate causes, accommodates short datasets, and, in principle, can be scaled efficiently to high dimensions using efficient Bayesian data assimilation algorithms. Crucially, it provides online tracking of causal roles, which may reverse intermittently, and facilitates a mathematically rigorous criterion for the causal influence range, revealing how far causal impacts propagate temporally. Its ability to dynamically detect causal influence ranges and handle high-dimensional systems by employing computationally efficient Bayesian techniques highlights its potential for applications in climate science, neuroscience, and other fields. Numerical tests on nonlinear systems with extreme events and regime transitions demonstrate the effectiveness and robustness of the framework.
It is worth highlighting how ACI differs from other state-of-the-art methods in terms of expressivity, computational complexity, and underlying assumptions. Different from most methods that identify cause-and-effect relationships by running the system forward and quantifying informational responses in the target variables’ entropy conditioned on past candidate causes, which is a predictive setting typical of information-flow-based methods, ACI instead formulates causality as an inverse problem in uncertainty quantification. Within the Bayesian data assimilation framework, candidate causal variables are inferred backward from observations of their potential effects. Thus, while many existing methods extrapolate causes forward in time, ACI traces causes back from observed effects, effectively interpolating rather than extrapolating causality. This distinction makes ACI more expressive in capturing instantaneous causal structures and, in practice, often more robust. From a computational standpoint, ACI can leverage well-developed Bayesian data assimilation tools that scale effectively to high-dimensional systems [evensen2003ensemble, law2015data]. In contrast, information transfer and other model-based methods often require computing the exact density functions and their derivatives, which is computationally demanding and subject to the curse of dimensionality. Finally, ACI requires only a single time series but assumes access to a dynamical model, which enables it to identify instantaneous causal relationships. In contrast, most time-series-based methods operate under model-free assumptions and therefore focus on aggregated or time-averaged causal links rather than on causality at specific moments. Because ACI incorporates model dynamics, it does not require long time series to detect causes and effects. Instead, it can infer instantaneous causal links from a single data point combined with model information. In contrast, most purely data-driven methods rely on long time series and can only reveal causal relationships in a time-averaged sense.
Future work will focus on several important directions. First, because model error is ubiquitous in practice, it is essential to study its impact on causal inference. Understanding how model error affects inferred causal relationships can create opportunities not only to detect and correct model error, but also to leverage this information to develop parsimonious models for complex dynamical systems. Similarly, observational data may contain additional noise that can affect causal inference. The ACI framework provides a natural way to analyze how observational noise contaminates causal detection. In such a context, data assimilation can serve as a pre-filtering step to reduce the influence of observational noise, which allows for a systematic examination of its role in shaping causal inference outcomes. Second, by building on the computationally efficient methods developed in Bayesian data assimilation, it is essential to apply ACI to higher-dimensional systems. Ensemble data assimilation is a well-established technique that routinely uses localization, covariance inflation, and other adjustments to overcome the curse of dimensionality [anderson2012localization, d2014scalable, anderson2007scalable]. As a result, the ensemble size (typically on the order of $O(100)$) is often chosen independently of the system’s dimension, which contributes to the scalability of Bayesian data assimilation and, consequently, of ACI. However, the sampling errors inherent in ensemble methods may introduce additional biases. Thus, it is necessary to examine whether such errors affect ACI. In addition, various coarse-grained approaches, such as approximating the full distribution with a Gaussian, are widely used in practice. Although these approximations introduce errors in computing the PDF, they are robust and highly accurate when computing diagnostic indicators, such as the relative entropy, that are used for model identification and sensitivity analysis. Thus, they may substantially accelerate ACI in practice. Third, building upon the current forward-in-time CIR analysis, the framework allows for the development of a crucial complement that focuses on dying the backward-in-time event attribution to past causes. The forward and backward CIR tools would support distinguishing different types of tipping points in Earth systems, identifying the causes of regime switching, and detecting early warning signals. Finally, continuous-in-time observations are adopted in this work for simplicity of presentation. Nevertheless, an analogous framework can be naturally extended to discrete-in-time observations using the corresponding Bayesian data assimilation algorithms. An interesting scientific question in this context is to understand the uncertainty in causal detection arising from the discrete nature of the data, particularly at time intervals between successive observations. This line of inquiry could also highlight an important distinction between ACI and purely data-driven methods. In ACI, the dynamical model provides a natural way to interpolate causal relationships at time instants when observations are unavailable. Therefore, it may help bridge observational gaps and maintain temporal continuity in causal inference.
@@
In Section 4.3.3 of the Supplementary Information (ACI on Real-World ENSO Data), daily and monthly observations for the state variables of the ENSO model were curated from the following reanalysis datasets: NCEP Global Ocean Data Assimilation System (GODAS) and NCEP-NCAR Reanalysis 1 Project. All other data in the simulations are model-generated.
Code Availability
The MATLAB code used in the analyses and to generate the figures in this work can be found in the following GitHub repository: https://github.com/marandmath/ACI_code. The codebase also includes implementations for additional experiments and real-world observations that can be used for ACI and CIR analysis in the ENSO case study.
@@
Supplementary Information
@@ A.2.1 Underlying Dynamics
We assume that the evolution of the partially observed process $(\mathbf{x},\mathbf{y})$ is governed by the following stochastic system of coupled Itô diffusions over $t\in[0,T]$ [rozovsky2012stochastic]:
@@
where $\mathbf{W}_{1}\in\mathbb{R}^{d_{1}}$ and $\mathbf{W}_{2}\in\mathbb{R}^{d_{2}}$ are two real-valued independent Wiener processes that are also independent from the distribution of $(\mathbf{x}(0),\mathbf{y}(0))$. We note that all subsequent random ordinary and stochastic differential equations in this document are written in differential form. In (1.1), we assume the following [rozovsky2012stochastic, ikeda2014stochastic, oksendal2003stochastic]. (I) Almost every sample path of $\mathbf{x}$ and $\mathbf{y}$ is continuous in $[0,T]$. (II) The functions appearing in (1.1) are predictable for each $\mathbf{x}\in\mathbb{R}^{k}$ and $\mathbf{y}\in\mathbb{R}^{l}$. (III) Sufficient regularity conditions are enforced such that (1.1) has a unique solution. (IV) The diffusion coefficients of $\mathbf{x}$ do not depend on $\mathbf{y}$: $\boldsymbol{\Sigma}^{\mathbf{x}}_{m}(t,\mathbf{x},\mathbf{y})=\boldsymbol{\Sigma}^{\mathbf{x}}_{m}(t,\mathbf{x})$, for $m=1,2$. (V) The sum of the row-based Gramians of the observational noise feedback matrices, $(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})(t,\mathbf{x}):=\boldsymbol{\Sigma}^{\mathbf{x}}_{1}(t,\mathbf{x})\boldsymbol{\Sigma}^{\mathbf{x}}_{1}(t,\mathbf{x})^{\mathtt{T}}+\boldsymbol{\Sigma}^{\mathbf{x}}_{2}(t,\mathbf{x})\boldsymbol{\Sigma}^{\mathbf{x}}_{2}(t,\mathbf{x})^{\mathtt{T}}$, is invertible for each $t\in[0,T]$ and $\mathbf{x}\in\mathbb{R}^{k}$.
(IV)–(V) function as identifiability conditions, since they ensure that the conditional distribution of $\mathbf{y}$ given the data of $\mathbf{x}$ contains all available information about $\mathbf{y}$ in (1.1) [rozovsky2012stochastic].